OpenStudy (anonymous):

how to make p(x)=x^3+2x^2-6x-4 as a product of two factors?

OpenStudy (anonymous):

u can use factor and remainder theorem to factorise it

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

first find the factors of 4 the factors of 4 are 1,2,4

OpenStudy (anonymous):

thus put x =+1 or -1 or +2 or -2 or +4 or -4 to see that p(x) becomes 0

OpenStudy (anonymous):

here for x=2 p(2) =(2)^3+2(2)^2-6(2)-4=8+8 -12-4=0 hence p(2) =0 thus x-2 is one of the factors

OpenStudy (anonymous):

now u can divide x^3+2x^2-6x-4 by x-2 to get the other factors or mere arrangement of terms will also give u the other factors as follows

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

p(x)=x^3+2x^2-6x-4 =x^2 -2x^2 +4x^2 -8x +2x = x^2 (x-2) +4x(x-2) +2(x-2) =(x-2)(x^2 +4x +2)

OpenStudy (jennychan12):

just factor. p(x) = x^3+2x^2-6x-4 p(x) = x^2(x+2)-2(x+2) p(x) = (x^2-2)(x+2)

OpenStudy (anonymous):

@matricked thanks, thats what i got (:

OpenStudy (anonymous):

welcome dear